OpenStudy (anonymous):

How does this problem work? I can't remember. Is it any different from the other ones? 2 • 3 - 2 • x = 2(3 - 11) And please don't leave negative remarks like "you should know this" I don't wanna hear it!

OpenStudy (radar):

The difference is to use the proper order of operations. You have parenthesis, subtraction, and multiplication. With those operations first get rid of the parenthesis then the multiplication, then the subtraction.

OpenStudy (radar):

\[2\times3 -2\times x=6 - 22\]\[6-2x =-16\]Now subtract 6 from both sides of the equation giving you: -2x = -22 Now divide both sides by -2 getting: x = 11. There was some algebraic operations in there, but I figure you know that part.

OpenStudy (radar):

There is some memory helpers. one is "My Dear Aunt Sally" Which shows an order of Multiplication, Division, Addition, and Subtraction. There are more involving exponents and to proceed from left to right.

OpenStudy (anonymous):

Thanks!!!

OpenStudy (radar):

PEDMAS is a more comphrensive nmenomic

OpenStudy (anonymous):

Ok thank you

OpenStudy (radar):

That would be, Parenthesis, Exponents, division, multiplication, addition and subtraction.

OpenStudy (anonymous):

Yes, i have heard of PEMDAS

OpenStudy (radar):

Good luck with your studies.

OpenStudy (anonymous):

Thank you!